"Micheal drove to a friends house at a rate of 40mph. He returned the same way at a rate of 45mph. The driving time for the whole trip was 4hours. What is the distance Micheal travled?"
Answer by checkley77(12569)

D=RT
D=40T
D=45(4-T)
40T=45(4-T)
40T=180-45T
85T=180
T=180/85
T=2.118 HOURS FOR THE FIRST TRIP.
4-2.118=1.882 HOURS FOR THE RETURN TRIP.
D=40*2.118=84.7 MILES.
D=45*1.882=84.7 MILES.
84.7*2=169.4 MILES IS THE DISTANCE TRAVELED.